the equation 1/2x(squared)+4x=0 has two solutions. whats the other?
Answer by priya(51)

the given equation is (1/2)x^2 +4x = 0.
first try to make the equation in a standard form like x^2 + x = 0

taking common denominator,
we get (x^2 + 8x)/2 = 0
x^2 + 8x = 0 ( 2*0 = 0,bring 2 on the right side by cross multiplication)
now take x as common,
x(x + 8) = 0.
therefore the values of x are 0 and x +8 =0 gives ,x = -8.
the answers are 0 and -8.
want to learn maths online..visit my site!!